# Explainable AI needs formal notions of explanation correctness

### v1

@@ 2 Purported purposes of XAI
Identification of intervention targets: It is frequently assumed that XAI could be used to identify features, the manipulation of which would change a model’s output, a task also known as algorithmic recourse  [21, 60]. For example [see 60], a bank might use an ML model to predict the return probability of a loan. For a known model and a given input, XAI would then be able to recommend changes of input variables (e.g., ‘increase salary’) to turn a negative outcome into a positive one. Similarly, it is assumed that XAI can help to verify that protected attributes (e.g., gender, race) do not influence model decisions. In an intensive care unit, an ML model might be used to predict mortality or other severe outcomes. Using XAI to identify possible intervention targets, such as medications, in this context [e.g., 6] implies that interventions have real-world consequences on the target variable beyond just changing the model output.
@@ 3 Existing feature attribution methods do not serve important purposes
Research has shown, though, that the assertions above do not hold in general for several of the most popular XAI methods. Thus, these methods fail to reliably serve the purposes mentioned in Section 2. We will demonstrate this next using two minimal examples.
@@ 3.4 Existing feature attribution methods violate common assertions
Another critical observation from Examples A and B is that $X_{2}$ cannot be independently manipulated but only jointly with $X_{1}$, due to its causal influence on $X_{1}$. Any causally feasible joint intervention on both variables would, however, not have a systematic effect on the model’s classification, and it would have zero effect on $Y$ in the real world. Thus, suppressor attributors also violate assertion A2 with the consequence that they cannot reliably suggest intervention targets – the manipulation of which would influence a model’s output or even the target variable. The insufficiency of counterfactual explanations to enable algorithmic recourse has also been noted in Karimi et al. [32].
@@ 4 Structural limitations of current XAI research
The results presented above have been established through joint theoretical analyses of data-generating processes, ML models, and feature attribution methods as well as through simulations using synthetic data with known ground-truth explanations. These techniques are not currently part of the standard toolkit for assessing the quality of explanations and XAI methods, pointing out the following fundamental structural limitations in the way the field assesses itself.
@@
The simplifying assumption of independent features is inherent to most XAI methods, be it explicit or implicit. Assuming independence is thereby in line with the common conception that the main benefit of multivariate modeling is to combine (independent) information in order to leverage non-linear interactions in the data. However, this perspective overlooks that an equally important task of multivariate models is to denoise interrelated features, which is achieved by removing task-irrelevant signals.
Incorrectly assuming independence can lead to all of the described misinterpretations. On the other hand, in rare cases where features can indeed be assumed to be independent, the same interpretations can often be shown to be valid.
@@
Various authors [e.g., 49] make a distinction between “explainable AI”, which would include feature attribution methods, and “interpretable AI”, which would include model architectures that are “understandable” by humans due to their simplicity. These are also referred to as “glassbox” models [46], and examples include linear models, models with sparse coefficients, and decision trees. However, what exact interpretations such models presumably afford is rarely stated. In the above examples A and B, linear models assign non-zero weights to suppressor variables, prohibiting certain desired interpretations and precluding certain actionable consequences that would be valid if the features were statistically independent. Thus, simple models share their reliance on the unrealistic assumption of feature independence with many feature attribution methods.
@@
Several studies [e.g., 26, 10] consider human judgments for XAI validation, where human experts either annotate inputs ex ante to provide ground-truth explanations or are asked to judge the quality of explanations ex post. Such approaches are insufficient as (sole) validations due to the possibility of both Type-I and Type-II errors in human judgments. For example, there may be complex statistical patterns in the data that are leveraged by ML models but (currently) unknown to humans. Similarly, human experts may hold wrong beliefs based on weak prior evidence. Human-in-the-loop studies also bear a certain danger of circular reinforcement of wrong beliefs, whereby humans may adapt their judgment to previously provided incorrect explanations.
@@ 5 Towards XAI for quality assurance
Based on formal criteria, it is possible to design ground-truth benchmark datasets that are realistic yet generated from a known and controlled parametric distribution. Various authors have proposed datasets in which the features sharing actual information with the prediction target are known by construction [28, 27, 70, 66, 5, 71, 15, 14, 44, 45]. This can be used to quantify specific definitions of explanation performance. In Oramas et al. [45], synthetic image datasets are introduced, where color manipulations of predefined object parts serve as class-related features used as ground-truth explanations. In Wilming et al. [66] and [15], a range of popular XAI methods in combination with distinct neural network architectures were benchmarked on linear and non-linear classification problems. In [44], structural magnetic resonance imaging (MRI) data were superimposed with synthetic brain lesions and the effect of pre-training on explanation performance in lesions classification tasks was studied. Wilming et al. [67] introduced a gender-balanced text dataset and associated gender classification tasks, which allows for quantifying explanation performance and biases in explanations. These datasets are publicly available.
@@ Acknowledgements
This work was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 758985) and the “Metrology for Artificial Intelligence in Medicine (M4AIM)” programme funded by the Federal Ministry for Economy and Climate Action (BMWK) in the frame of the QI-Digital Initiative.

### v2

@@ 2 Purported purposes of XAI
Identification of intervention targets: It is frequently assumed that XAI could be used to identify features, the manipulation of which would change a model’s output, a task also known as algorithmic recourse  [60]. For example [see 60], a bank might use an ML model to predict the return probability of a loan. For a known model and a given input, XAI would then be able to recommend changes of input variables (e.g., ‘increase salary’) to turn a negative outcome into a positive one. Similarly, it is assumed that XAI can help to verify that protected attributes (e.g., gender, race) do not influence model decisions. In an intensive care unit, an ML model might be used to predict mortality or other severe outcomes. Using XAI to identify possible intervention targets, such as medications, in this context [e.g., 6] implies that interventions have real-world consequences on the target variable beyond just changing the model output.
@@ 3 Existing feature attribution methods do not serve important purposes
Research has shown, though, that the assertions listed above do not hold in general for several of the most popular XAI methods. Thus, these methods fail to reliably serve the purposes mentioned in Section 2. We will demonstrate this next using two minimal examples.
@@ 3.4 Existing feature attribution methods violate common assertions
Another critical observation from Example B is that $X_{2}$ cannot be independently manipulated but only jointly with $X_{1}$, due to its causal influence on $X_{1}$. Similarly, $X_{1}$ and $X_{2}$ are coupled through the common noise ${\mathbf{H}}$ in Example A. Depending on the nature of $X_{2}$, it then might or might not be possible to intervene at $X_{2}$ alone by setting its value independent of $H$. In both examples, any causally feasible intervention (through $X_{2}$ or $H$) would have no effect on $Y$ in the real world. In Example B, it would not even affect the model output, as the model is invariant to changes in $X_{2}$ by construction. In Example B, possible interventions on $X_{2}$ could affect the model output; however, not in ways that would correlate with changes in $Y$. Such interventions would create out-of-distribution data that do not follow the data generating process (similar to pertubations based faithfulness metrics discussed below). Since the model has not seen data perturbed this way during training, a discussion of the effect of such pertubations on the model output does not appear to be particularly useful with respect to understanding the model.
Together, both examples illustrate that suppressor attributors also violate assertion A2 with the consequence that they cannot reliably suggest feasible intervention targets. The insufficiency of counterfactual explanations to enable algorithmic recourse has been laid out in detail in Karimi et al. [32].
@@ 4 Structural limitations of current XAI research
The results presented above have been established through joint theoretical analyses of data-generating processes, ML models, and feature attribution methods as well as through simulations using synthetic data with known ground-truth explanations. These techniques are not currently part of the standard toolkit for assessing the quality of explanations and XAI methods, pointing to the following fundamental structural limitations in the way the field assesses itself.
@@
Most XAI methods explicit or implicit assume independent features. This is in line with the common conception that the main mechanism by which multivariate models achieve their predictive power is to combine (independent) information in order to leverage non-linear interactions in the data. However, this perspective overlooks that an equally important task of multivariate models is to denoise interrelated features, which is achieved by removing task-irrelevant signals. Incorrectly assuming independence can lead to violations of assertion A1 and assertion A2, and, thereby, to all of the described misinterpretations. On the other hand, in cases where features can indeed be assumed to be independent, the same interpretations can often be shown to be valid. [32], for example, show that correct algorithmic recourse is possible under feature independence.
@@
Various authors [e.g., 49] make a distinction between “explainable AI”, which would include post-hoc feature attribution methods, and “interpretable AI”, which would include model architectures that are “understandable” to humans due to their simplicity. These are also referred to as “glassbox” models [46], and examples include linear models, models with sparse coefficients, and decision trees. However, what exact interpretations such models presumably afford is rarely stated. In the above examples A and B, linear models assign non-zero weights to suppressor variables, prohibiting certain desired interpretations and precluding certain actionable consequences that would be valid if the features were statistically independent. Thus, simple models can share unrealistic assumptions such as the assumption on independent features with many feature attribution methods.
@@
Several studies [e.g., 26, 10] consider human judgments for XAI validation, where human experts either annotate inputs ex ante to provide ground-truth explanations or are asked to judge the quality of explanations ex post. While important, such approaches are insufficient as (sole) validations due to the possibility of both Type-I and Type-II errors in human judgments. For example, there may be complex statistical patterns in the data that are leveraged by ML models but (currently) unknown to humans. Similarly, human experts may hold wrong beliefs based on weak prior evidence. Human-in-the-loop studies also bear a certain danger of circular reinforcement of wrong beliefs, whereby humans may adapt their judgment to previously provided incorrect explanations.
@@ 5 Towards XAI for quality assurance
Based on formal criteria, it is possible to design ground-truth benchmark datasets that are realistic yet generated from a known and controlled parametric distribution. Various authors have proposed datasets in which the features sharing actual information with the prediction target are known by construction [28, 27, 66, 5, 71, 15, 14, 44, 45]. This can be used to quantify specific definitions of explanation performance. In Oramas et al. [45], synthetic image datasets are introduced, where color manipulations of predefined object parts serve as class-related features used as ground-truth explanations. In Wilming et al. [66] and [15], a range of popular XAI methods in combination with distinct neural network architectures were benchmarked on linear and non-linear classification problems. In [44], structural magnetic resonance imaging (MRI) data were superimposed with synthetic brain lesions and the effect of pre-training on explanation performance in lesions classification tasks was studied. Wilming et al. [67] introduced a gender-balanced text dataset and associated gender classification tasks, which allows for quantifying explanation performance and biases in explanations. These datasets are publicly available.
@@ Acknowledgements
This work was supported by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant agreement No. 758985), the European Partnership on Metrology (Grant agreement 22HLT05), and the Metrology for Artificial Intelligence in Medicine (M4AIM) programme funded by the Federal Ministry for Economy and Climate Action (BMWK) in the frame of the QI-Digital Initiative.